What are Stock Index Futures? How do they differ from Stock Index Options.? Give a detailed example of how a trader might find it a better idea to hedge with Stock Index Futures rather than with Stock Index Options.What are Stock Index Futures? How do they differ from Stock Index Options.? Give a detailed example of how a trader might find it a better idea to hedge with Stock Index Futures rather than with Stock Index Options.
Answer #1

Stock Index Futures refers to future derivative instruments for a stock market index. For example: Dow Jones Futures. Stock market index represents the state of all its components stocks.

Stock index options are option derivative instruments which have different strike price for call/put, option premium values.

  • In Stock Index Futures there is no risk of theta decay which is option losing values closer to maturity
  • There is higher margin requirement to take a position in futures than in option buying.
  • Index futures are however cheaper as compared to index options selling which requires huge margin.
  • There is no option greeks involved in index futures.
  • There are multiple option strategies available as compared to futures.
  • Index futures have mark to market gain/loss at end of each trading day as against stock index options in case of carry forward positions.

Example showing hedging with Index futures:

Stock: Amazon (AMZN)

Stock index: Nasdaq 100 Index (containing AMZN as its component)

Existing position: The trader is short on AMZN Stock futures having beta of 1.2 (assumed) to the value of $100,000. Suppose the value of AMZN is $100 ( for the sake of brevity). Lot size = 100 Now, no. of futures bought 100000/(100*100)=10.

Hedging: Index always have beta of 1. Now if the AMZN stock to move up, he will be incurring losses. To hedge the same, the trader initiates the following position in Index Futures Long ( Counter position to existing position). Index Lot size =100. Future price =200

=1.2*100*100*10/100*200 = 6 Index Futures Long.

Under such circumstances, if the AMZN stocks moves up, the trader would be beneitted by increase in the value of index futures to the same value and his overall value of profit/loss = 0 (Zero). This is perfect hedge strategy using index futures. Similarly, a trader can use partial too.

In Index options, the trader cannot adequately determine hedge ratio. Further, the involvement of option greeks of index options make index options hedging a complex strategy for hedging against stock futures